OpenStudy (anonymous):

Draw an 8 by 5 Fibonacci rectangle and show the next Fibonacci rectangle after removing a square gnomon

OpenStudy (phi):

As you know (?) the Fibonacci numbers are created by starting with the sequence 1,1 you add the last two numbers in the sequence to get the next one, so 1+1=2 is the next. 1,1,2 now add 1+2=3 1,1,2,3,5,8,13 and so on. The next after 13 is 8+13= 21 Draw an 8 by 5 Fibonacci rectangle means draw a rectangle with sides of 5 and 8 (which are Fibonacci numbers)